OpenStudy (aravindg):

1.for non zero real nos: x and y the equality [x+y]^2 sec^2\[\theta\]=4xy is possible only when a)x+y=1 b)x+y=-1 c)x-y=0 d)x+y=0

OpenStudy (amistre64):

for non zero real nos: x and y the equality [x+y]^2 sec^2(θ) = 4xy, hmmm

OpenStudy (amistre64):

we could back pedal into it to see maybe?

OpenStudy (aravindg):

???

OpenStudy (amistre64):

what is theta, anything?

OpenStudy (aravindg):

angle

OpenStudy (amistre64):

good, cause i was thinking it might be been soup for a minute :)

OpenStudy (aravindg):

answer?

OpenStudy (amistre64):

i dunno an answer; but is sounds plausible that when x and y are opposites that you get a few zeros

OpenStudy (amistre64):

then again ....

OpenStudy (amistre64):

y = x; [x-x]^2 sec^2(θ) = 4x^2 not right

OpenStudy (amistre64):

myin might be a good bet for doing these number bending stuffs

OpenStudy (aravindg):

???

OpenStudy (aravindg):

i neeed the answer

OpenStudy (amistre64):

openanswer.com is 3 doors down; this is more of a study group in my eyes

OpenStudy (amistre64):

at any rate, i got no answers for it, so good luck with it ;)

OpenStudy (aravindg):

helppppp

OpenStudy (anonymous):

try wolfram alpha

OpenStudy (anonymous):

(not a user; a site)

OpenStudy (aravindg):

helpp

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

from what topic or book does this question come?

OpenStudy (amistre64):

im assuming its the help me now i need an answer book; just a guess tho

OpenStudy (anonymous):

maybe it is the book of soup (none for you though)

myininaya (myininaya):

so it does not have to hold for all real x right? it holds for some x that i see if x=y, then we have (2x)^2sec^2(theta)=4x^2 4x^2sec^2(theta)=4x^2 and we know sec^2(theta) can be 1 sometimes since cos^2(theta) can be 1

OpenStudy (anonymous):

we can try \[x=y\] right?

OpenStudy (anonymous):

oh i am a little slow i see. hello myininaya

OpenStudy (anonymous):

yes i think for some values of theta x = y will work

myininaya (myininaya):

hey satellite i'm tired i think i'm gonna eat and watch tv

OpenStudy (anonymous):

oh you wrote that out too. ok done. since it says "only possible if" then you must be right. is antiques roadshow on?

OpenStudy (anonymous):

now the question is, why not for any other value...

OpenStudy (aravindg):

:(

OpenStudy (aravindg):

:(

OpenStudy (aravindg):

guys

OpenStudy (aravindg):

:(

OpenStudy (aravindg):

u cant answer?????

OpenStudy (aravindg):

helpppp

myininaya (myininaya):

i think we agreed to x=y